OpenStudy (anonymous):

find an equationof the line containing the given pair of points. (-1,-2), (-9,-8) the equation ofmthe line in slope intercept form is y=?

OpenStudy (anonymous):

@acxbox22

OpenStudy (anonymous):

is this y1-y2=m(x1-x2)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

I'm gonna try to work it out before I give you the answer... two seconds

OpenStudy (acxbox22):

y=3/4x-5/4 is the answer

OpenStudy (anonymous):

@acxbox22 I couldn't figure it out, would you mind showing me how you did it?

OpenStudy (anonymous):

@acxbox22 I used y1-y2=m(x1-x2) but I didnt get that answer

OpenStudy (acxbox22):

that explains why.. the equation is y-y1=m(x-x1) where y1 and x1 are the x and y of either two points and m is slope

OpenStudy (anonymous):

ohh thanks